Operation Summary for August 19, 2011 - August 22, 2011 |
The Integrated Luminosity for the period from 8/15/11 to 8/22/11 was 31.61 pb-1 (inverse picobarns). NuMI reported receiving 6.63E18 protons on target during this same period. Friday August 19 The day shift began with Operations monitoring the Tevatron (TEV) store 8998 (74.86E30) and the Recycler stash (20.0E10), with the Antiproton Source (Pbar) stacking (143.1E10), with the Main Injector (MI) operating normally, but with the Pelletron off, with Linac and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations terminated store 8998 at 10:16 AM. Operations loaded the TEV with one low intensity bunch of protons at 10:59 AM for a Proton only store. At 12:59 PM, operators reported bad Booster extraction positions and beam permit trips; then a Booster kicker (B:MKS05) power supply tripped off and would not reset; Booster personnel investigated. A Booster RF person began repairs on a RF station (BRF5). All Booster work was completed by 2:36 PM and Operations established beam to NuMI, MiniBooNE, and Pbar. Operators reported at 3:15 PM that BRF5 tripped off for the first of three times in five minutes; a RF person investigated. During the evening shift, Operations established the 3x3 study store at 4:07 PM. Operations established beam to the Meson FTBF T-1015 experiment around 4:45 PM. At 5:40 PM, the 3x3 store aborted with no quench. Operators put the TEV through a wet squeeze at 6:35 PM. Operations put the TEV into shot setup at 7:40 PM. Operations disabled beam to T-1015 at 8 PM. Operations established store 9004 at 8:46 PM with an average initial luminosity (AIL) of 115.9E30. CHL Cryo System personnel reported at 9:52 PM that the cold compressor at TEV sector DA tripped off; they switched D0 to run off of the TEV header. Saturday August 20 The midnight shift began with Operations monitoring store 9004 (66.63E30), the stash (170.3E10) and the Pbar stack (18.4E10), and with the NuMI and MiniBooNE taking beam. During the midnight shift, MI personnel reported that they had completed the Pelletron work; operators had the Pelletron and enclosure MI-31 searched and secured by 12:42 AM. Operations terminated store 9004 at 3:50 AM, put the TEV into shot setup, and established store 9005 at 5:13 AM with an AIL of 115.05E30. Operators reported a short downtime due to problems with a Linac RF station (LRF2). During the day shift, Operations established beam to T-1015 at 10:10 AM. At 10:18 AM, the TEV quenched at A4, B1, BC, and BD due to a lightning strike. Operators prepared D0 for an access. At 12:55 PM, D0 came out of access and CDF went into access. MI personnel reported at 2:02 PM that the Pelletron conditioning had gone well, but then suffered a full discharge; MI personnel reported that the third level correctors had failed. Operators prepared the Pelletron for access. At 3:04 PM, the TEV sector F4 wet engine tripped off; Cryo System personnel began repairs. At 3:23 PM, BRF5 began sparking constantly; operators bypassed the station and brought on BRF19 in its place. CDF came out of access at 3:23 PM. During the evening shift, Cryo System personnel reported at 4:35 PM that the F4 wet engine and the cryo system were back online. Operators began ramping the TEV at 5:10 PM. Operators put the TEV through a dry squeeze at 5:28 PM. Operations put the TEV into shot setup at 7:19 PM. Operations disabled beam to T-1015. Operations established store 9009 at 8:22 PM with an AIL of 136.36E30. Sunday August 21 The midnight shift began with Operations monitoring store 9009 and the stash, with Pbar stacking, and with the NuMI and MiniBooNE taking beam. During the midnight shift, operators reported at 12:28 AM that a Recycler kicker (R:KPS3A) power supply had tripped off on a temperature/Fluorinert fault indication and would not reset; operators contacted personnel to fill the kicker, which required an access for the Pelletron to be off. BRF16 tripped off at 2:11 AM and wouldn’t reset; operators bypassed the station and brought up BRF5 in its place. Operations terminated store 9009 at 3:34 AM. At 3:25 AM, the BRF13 power supply tripped off on a ground fault indication; operators bypassed the station and brought BRF16 online. Operations put the TEV into shot setup at 3:46 AM and established store 9010 at 4:50 AM with an AIL of 125.19E30. MI personnel accessed to fill R:KPS3A with Fluorinert at 5:36 AM. Operators reported various Booster problems to a Booster person at 6:47 AM that were holding off beam; the Booster person recommended an RF person look at BRF13. A Booster RF person came in and had BRF13 back online and beam from Booster by 6:55 AM. During the day shift, Operations terminated store 9010 at 1:03 PM and prepared D0 for an hour access. Operations established beam to T-1015 at 2:01 PM. Operators put the TEV through a dry squeeze at 2:37 PM and then a wet squeeze at 3:52 PM. During the evening shift, Operations put the TEV into shot setup at 4:16 PM and established store 9014 at 5:32 PM with an AIL of 108.5E30; operators reported that a sextupole magnet power supply (C:S5F3) tripped off during the low beta squeeze, which was why the luminosity was low. At 5:51 PM, the TEV quenched at TEV sector A1; operators reported that the A17 separator vacuum pumps had tripped off. A TEV person came in to investigate. Operations disabled beam to T-1015 at 7:10 PM. The TEV person reported at 7:49 PM that he was closing a vacuum valve (BV7M) to help determine where the vacuum leak was. Operators searched and secured the Pelletron around 8 PM and put it into conditioning mode. Vacuum personnel reported at 10:15 PM that they found and temporarily repaired a vacuum leak on the TEV A17 bellows; the repair will take time to cure; no store. Monday August 22 The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with the NuMI and MiniBooNE taking beam. During the midnight shift, a Pelletron person reported at 12:03 AM that the Pelletron was back in service. Operators reported at 6:01 AM that the Pelletron bias voltage (R:BIASV) kept tripping off; a Pelletron person had operators restore the Electron Cooling correctors. At 6:24 AM, operators reported more problems with the Pelletron and called in people to help. Operators reported short downtimes due to LRF1, LRF3, LRF5, a Klystron RF station (KRF1), BRF7, and BRF9. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow an 8 GeV coalescing study, TEV F4 cryo system compressor work, and a TEV 3x3 study store later this evening. |
|
More Information Comments and Suggestions |
| last modified 08/22/2011 email Fermilab |
| Security, Privacy, Legal | |